In how many different ways can you answer a multiple choice test that has 9 questions and 3 choices for each answer
Answer by stanbon(75887)

In how many different ways can you answer a multiple choice test that has 9 questions and 3 choices for each answer
-------------
1 questin----3 ways
2 questions--3*3 ways
3questions---3*3*4 = 3^3 ways
---
----
9 questions: 3^9 = 19683 ways
==================================
Cheers,
Stan H.
============